# Conservation-informed Graph Learning for Spatiotemporal Dynamics Prediction

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025conference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canada.booktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadaisbn: 979-8-4007-1245-6/25/08doi: 10.1145/XXXXXX.XXXXXXccs: Computing methodologies Modeling and simulationccs: Computing methodologies Neural networksccs: Applied computing Physics
@@ 2. Related Works
As one of the most popular framework, the seminal development of the physics-informed neural network (PINN) (Raissi et al., 2019; Jagtap et al., 2020) has enabled learning in small data regimes for forward and inverse analyses of PDEs, where the loss function is regularized by the PDE residual. Such a differentiable paradigm has kindled enthusiastic attention in the past few years, which, together with its variant schemes, has been demonstrated effective in modeling and discovering a wide range of complex systems including fluid flows (Raissi et al., 2020), subsurface transport (He and Tartakovsky, 2021), cardiovascular systems (Arzani et al., 2021), solid continua’s deformation (Niu et al., 2023), material constitutive relations (Rezaei et al., 2024), governing equation discovery (Chen et al., 2021), data superresolution (Ren et al., 2023), elastic waves (Rao et al., 2021), structural metamodeling (Zhang et al., 2020b, c), among many others. However, PINN generally has the need of explicit equations and the issue of scalability, resorting to the use of fully-connected neural networks and mesh-free point-wise training.
@@ 4. Methodology
We develop a learning scheme that conforms to the conservation law shown in Eq. 1. The schematic overview of our proposed CiGNN is shown in Figure 1 which illustrates an end-to-end learning pipeline of “Encoder-Processor-Decoder” framework.
@@ Acknowledgements.
The work is supported by the National Natural Science Foundation of China (No. 62276269, No. 92270118), the Beijing Natural Science Foundation (No. 1232009), and the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0620103). H. Sun and Y. Liu would like to thank the support from the Fundamental Research Funds for the Central Universities (No. 202230265 and No. E2EG2202X2). P. Ren would like to disclose that he was involved in this work when he was at Northeastern University, who has not been supported by Huawei Technologies.

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2025conference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709244isbn: 979-8-4007-1245-6/25/08ccs: Computing methodologies Modeling and simulationccs: Computing methodologies Neural networksccs: Applied computing Physics
@@ 2. Related Works
As one of the most popular framework, the seminal development of the physics-informed neural network (PINN) (Raissi et al., 2019; Jagtap et al., 2020) has enabled learning in small data regimes for forward and inverse analyses of PDEs, where the loss function is regularized by the PDE residual. Such a differentiable paradigm has kindled enthusiastic attention in the past few years, which, together with its variant schemes, has been demonstrated effective in modeling and discovering a wide range of complex systems including fluid flows (Raissi et al., 2020), cardiovascular systems (Arzani et al., 2021), solid continua’s deformation (Niu et al., 2023), material constitutive relations (Rezaei et al., 2024), governing equation discovery (Chen et al., 2021), data superresolution (Ren et al., 2023), elastic waves (Rao et al., 2021; Ren et al., 2024), structural metamodeling (Zhang et al., 2020b, c), among many others. However, PINN generally has the need of explicit equations and the issue of scalability, resorting to the use of fully-connected neural networks and mesh-free point-wise training.
@@ 4. Methodology
We develop a learning scheme that conforms to the conservation law shown in Eq. 1. The schematic overview of our proposed CiGNN is shown in Figure 1 which illustrates an end-to-end learning pipeline of “Encoder-Processor-Decoder” framework. We have released the source code and data at https://github.com/intell-sci-comput/CiGNN.
@@ Acknowledgements.
The work is supported by the National Natural Science Foundation of China (No. 62276269, No. 92270118), the Beijing Natural Science Foundation (No. 1232009), and the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0620103). H. Sun and Y. Liu would like to thank the support from the Fundamental Research Funds for the Central Universities (No. 202230265, No. E2EG2202X2). P. Ren would like to disclose that he was involved in this work when he was at Northeastern University, who has not been supported by Huawei Technologies.